Suppose a random variable is normally distributed. The probabilities for a85 and 142 are 10% and 65%, respectively. Find the mean and standard deviation to the nearest whole number.

1
Expert's answer
2021-12-07T11:19:28-0500

Mean=xf(x)

=(85*0.1)+(142*0.65)

=100.8

≈101

Variance="\sum (x-\mu)^2f(x)"

=(85-100.8)².01+(142-100.8)².0.65

=1128.3

sd=33.59